# Topological chiral superconductivity

### v1

@@ Abstract
We investigate a mechanism to produce superconductivity by strong purely-repulsive interactions, say in multi-layer graphene. The resulting superconductors break both time-reversal and reflection symmetries in the orbital motion of electrons, and exhibit non-trivial topological order. Our findings suggest that this topological chiral superconductivity is more likely to emerge near or between fully spin-valley polarized metallic phase and Wigner crystal phase. Unlike conventional BCS superconductors in fully spin-valley polarized metals, these topological chiral superconductors are only partially spin-valley polarized, even though the neighboring normal state remains fully spin-valley polarized. The ratios of electron densities associated with different spin-valley quantum numbers are quantized as simple rational numbers. Furthermore, many of these topological chiral superconductors exhibit charge-4 or higher condensation. One of the topological chiral superconductors is in the same phase as the “spin”-triplet $p+\textrm{i}p$ BCS superconductor, while others are in different phases than any BCS superconductors. The same mechanism is also used to produce anyon superconductivity between fractional anomalous quantum Hall states in the presence of a periodic potential.
@@ I Introduction
After the discovery of superconductivity in 1911 [1], the standard BCS mechanism for superconductivity was developed in 1957 [2], based on the electron pairing instability of Fermi liquid, caused by an effective attraction between electrons. In this paper, we explore a very different mechanism of superconductivity, which is caused by strong purely-repulsive interaction. The superconductivity from our mechanism is very different from BCS superconductivity.
In fact, the mechanism based on the charged anyons in chiral spin liquid [3, 4, 5], and related models [6, 7, 8, 9, 10, 11, 12], belongs to this class of mechanism (i.e. driven by purely-repulsive interactions). In this paper, we obtain superconductivity directly from electrons with repulsive interaction, without going through charged anyons and the associated anyon superconductivity. The resulting superconductivity may not be associated with electron pairing; charge-$4$ (and higher) condensation is also possible [11]. As a result, the resulting superconductivity usually carries non-trivial topological order [13, 14], which will be referred to as topological chiral superconductivity.
@@
Certainly, when interaction is weak, electrons will form a Fermi liquid. However, when interaction is strong, the electron motions are highly correlated. Since electron hopping is complex, the two electrons exchange their place via correlated motion, and the phase factor can be arbitrary. In this case, electrons may forget their Fermi statistics. Thus, electrons may form a superconducting state via the mechanism of anyon or boson superconductivity. The above idea is very rough, but may point to a right direction [12].
@@
In this paper, we discuss a concrete realization of the above idea in 2-dimensional space. We argue that a strong repulsive interaction may cause a chiral superconductivity that breaks time reversal and space reflection symmetry in orbital motion of electrons. Time-reversal symmetry breaking and/or spin-orbital coupling in orbital motion may further help this chiral superconductivity.
We will use the theory of anyon superconductivity developed in Ref. 8, 12. Since a fermion is a special case of anyon, the theory of anyon superconductivity applies to fermion superconductivity without change. We apply our theory to multi-layer graphene. We find many topological chiral superconductors driven by purely-repulsive interaction, whose ground-state energies may be lower than those of fully spin-valley polarized Fermi liquid (referred to as quarter Fermi liquid) and Wigner crystal at low densities.
@@
Quarter Fermi liquids and Wigner crystals have both been observed in experiments at low densities where the interaction effect is strong[15]. Furthermore, a strong coupling superconducting state that breaks time reversal and reflection symmetries in electron orbital motion was observed in Ref. 15 between quarter Fermi liquid and Wigner crystal, in tetralayer rhombohedral-stacked graphene without Moire pattern. Other superconductivities were also observed in bilayer [16, 17, 18, 19], trilayer [20, 21], and Moire-tetralayer [22] rhombohedral-stacked graphene systems. Some BCS-pairing mechanisms for those superconducting states were explored in Ref. 23, 24, 25, 26, 27, 28, 29, 30.
If a superconducting state is observed close to a quarter Fermi liquid, as in Ref. 15, one might expect the superconducting state to be also fully spin-valley polarized. In this case, we need an effective attraction between electrons to induce the pairing instability of the Fermi liquid.
@@
In this paper, we take a very different approach of using only strong repulsive interaction, which leads to very different chiral superconductors. The resulting chiral superconductors are not fully spin-valley polarized and are not induced by paring instability of Fermi liquid. Some of them are half spin-valley polarized (with two spin-valley components present and at equal density) while others are spin-valley un-polarized (with all four spin-valley components at equal density). We stress that, in our chiral superconductors, the ratio of different species of electrons are quantized as simple rational numbers. Thus, in contrast to BCS superconductors of quarter Fermi liquid, the transition from spin-valley partially-polarized chiral superconductors to quarter Fermi liquid cannot be continuous at zero temperature in the clean limit.
@@
Also, as we lower the electron density, a topological chiral superconductor is likely to change into a Wigner crystal, also via a first order transition. Thus topological chiral superconductivity is more likely to appear near the transition between quarter Fermi liquid and Wigner crystal, since all those phases are driven by strong repulsive interactions. As a strong coupling superconductor, the coherence length of a chiral superconductor is about the same as the inter-electron separation.
All these chiral superconductors are topological (i.e. carry non-trivial topological order); many of them have charge-4 or higher condensation and will break time reversal and space reflection symmetry in addition to carrying gapless chiral edge modes.
@@
Our calculation is not reliable enough to predict if a chiral superconductor can appear or not (i.e. can have energy below both quarter Fermi liquid and Wigner crystal). However, if a chiral superconductor (i.e. a superconductor that breaks time reversal and reflection symmetry) is observed in experiments near quarter Fermi liquid and Wigner crystal as in Ref. 15, our calculation suggests that it may be a partially spin-valley polarized topological superconductor, with properties described above.
@@
The calculated ground-state energies of these partially spin-valley polarized superconductors are lower, but still close to Hartree-Fock energy of a quarter Fermi liquid. To achieve a lower energy state than the quarter Fermi liquid, additional time-reversal symmetry breaking and/or spin-orbit interaction may be helpful. The strong geometric phase curvature near the bottom of the graphene band also contributes to breaking time-reversal symmetry through partial spin-valley polarization.
@@
We remark that, although most topological chiral superconductors belong to different phases than BCS superconductors, one of the topological chiral superconductors (which we denote as the $K_{2a}$-chiral superconductor) belong to the same phase as the “spin”-triplet $p+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp$ BCS superconductor. Here “spin” corresponds to a pair of spin-valley quantum numbers, and may not be the electron spin. Such a $K_{2a}$-chiral superconductor can be induced by a purely-repulsive interaction. It can also be induced by a pairing instability of a half Fermi liquid caused by an effective attractive interaction. Here, half Fermi liquid refers to a Fermi liquid formed by two species of electrons of equal density. We note that very recently, a superconducting state has been observed next to a half Fermi liquid in rhombohedral trilayer graphene in Ref. 21. The $K_{2a}$-chiral superconductor also has similarly low energy to other chiral superconductors at lower densities. Thus, it may also appear near the quarter Fermi liquid and Wigner crystal, as in rhombohedral tetralayer graphene in Ref. 15.
@@
II Chiral superconductivity driven by purely-repulsive interactions
@@
We now examine the energetics of this state. Let $N_{\uparrow}$ ($N_{\downarrow}$) be the total number of spin-$\uparrow$ (spin-$\uparrow$) electrons. The total angular momentum of the above state is
@@
We will use (II) as a trial wave function for the associated chiral superconductor. Such a wave function is determined by $\bar{K}^{\pm}$, which must satisfy some conditions, as discussed in detail in Appendix. In the main text, we will just summarize the results. First,
@@
$=\frac{1}{2}\sum_{I,J}N_{I}K_{IJ}N_{J}-\frac{1}{2}\sum_{I}N_{I}K_{II}$ (20)
where $N_{I}$ is the number of species-$I$ electrons. In order to describe a superconducting or superfluid state, $K$ is required to have a single zero eigenvalue so that the total angular momentum does not contain the $N^{2}$ term:
@@
The corresponding eigenvector $f_{I}$ describes a gapless mode within the otherwise gapped chiral superconducting state. In order to be a superconducting state, we also require the electron density fluctuations giving rise to this mode to be net positive; otherwise, it would describe a charge neutral superfluid mode instead of a superconductor. $f_{I}$ is proportional to the density of species-$I$ electrons. Thus, we also require
@@
If the electron dispersion has a form $\varepsilon=c_{2}k^{2}+c_{4}k^{4}$, then the kinetic energy per electron can be expressed as
@@
where $n_{e}$ is the total density of the electrons. We find (see Appendix):
$Z_{2}$ $=\sum_{IJ}f_{I}\bar{K}_{IJ}f_{J}+\frac{2g_{2}}{\pi n_{e}},$
$Z_{4}$ $=\sum_{I}2f_{I}(\sum_{J}f_{J}\bar{K}_{IJ})^{2}+\frac{16g_{4}}{\pi^{2}n_{e}^{2}},$
$\bar{K}$ $\equiv\bar{K}^{+}_{IJ}+\bar{K}^{+}_{IJ}.$ (26)
where the dimensionless ratios, $\frac{g_{2}}{n_{e}}$ and $\frac{g_{4}}{n_{e}^{2}}$, can be computed numerically via Monte Carlo method, with a current error about 10%.
If we assume the electrons interact via Coulomb interaction $\frac{e^{2}}{\epsilon r}$, then the interaction energy per electron is given by
@@
and $g_{IJ}(z)$ is the electron pair distribution function which needs to be computed numerically. We find the following approximate fitting for $V_{IJ}$
@@
with error $\sim 0.03$ (see Appendix), where $\Theta(0)=0$ and $\Theta(x>0)=1$. When $\bar{K}_{IJ}=K_{\text{max}}$, the energy $E_{\text{kin}}$ and $E_{\text{int}}$ happen to be the same as the one component case $\bar{K}=(K_{\text{max}})$. We will use this property to compute $E_{\text{kin}}$ and $E_{\text{int}}$ for the case $\bar{K}_{IJ}=K_{\text{max}}$.
There are many $K$-matrices that satisfy (21) and (22). To determine which $K$-matrices give rise more stable chiral superconductors, we compute $Z_{2}$, $Z_{4}$, and $V$. The total energy per electron is given by
@@
To estimate $\bar{n}_{e}$ for the tetralayer graphene at $\gamma=4$, we use the band dispersion $10$meV $=c_{4}(0.1/a_{0})^{4}$ where $a_{0}=0.246$nm $=4.65a_{B}$, giving $\bar{n}_{e}=8.5\times 10^{12}$ cm-2. To estimate $\bar{n}_{e}$ for the tetralayer graphene at $\gamma=2$, we use band dispertion $30$meV $=c_{2}(0.1/a_{0})^{2}$, giving $\bar{n}_{e}=9.8\times 10^{12}$ cm-2.
@@
Because the densities in different spin-valley components have quantized ratios and is very different between those chiral superconducting states and quarter Fermi liquid, the zero-temperature transition between chiral superconducting states and quarter Fermi liquid is first order. This is a key prediction of this paper.

### v3

@@ Abstract
We investigate a mechanism to produce superconductivity by strong purely repulsive interactions for flat dispersion $\varepsilon\sim k^{4}$, without using pairing instability in Fermi-liquid. The resulting superconductors break both time-reversal and reflection symmetries in the orbital motion of electrons, and exhibit non-trivial topological order. Our findings suggest that this topological chiral superconductivity is more likely to emerge near or between fully spin-valley polarized metallic phase and Wigner crystal phase. These topological chiral superconductors can be fully or partially spin-valley polarized. For partial spin-valley polarization, the ratios of electron densities associated with different spin-valley quantum numbers are quantized as simple rational numbers. Furthermore, many of these topological chiral superconductors exhibit charge-4 or higher condensation, neutral quasiparticles with fractional statistics, and/or gapless chiral edge states. Two of the topological chiral superconductors are in the same phases as the “spin”-triplet or spinless $p+\textrm{i}p$ BCS superconductor, while others are in different phases than any BCS superconductors. The same mechanism is also used to produce anyon superconductivity between fractional anomalous quantum Hall states in the presence of a periodic potential.
@@ I Introduction
After the discovery of superconductivity in 1911 [1], the standard BCS mechanism for superconductivity was developed in 1957 [2], based on the electron pairing instability of Fermi liquid, caused by an effective attraction between electrons. In this paper, we explore a very different mechanism of superconductivity, which is caused by strong purely repulsive interaction. The superconductivity from our mechanism is very different from BCS superconductivity.
In fact, the mechanism based on the charged anyons in chiral spin liquid [3, 4, 5], and related models [6, 7, 8, 9, 10, 11, 12], belongs to this class of mechanism (i.e. driven by purely repulsive interactions). In this paper, we obtain superconductivity directly from electrons with repulsive interaction, without going through charged anyons and the associated anyon superconductivity. The resulting superconductivity may not be associated with electron pairing; charge-$4$ (and higher) condensation is also possible [11]. As a result, the resulting superconductivity usually carries non-trivial topological order [13, 14], which will be referred to as topological chiral superconductivity.
@@
Certainly, when interaction is weak, electrons will form a Fermi liquid. However, when interaction is strong, the electron motions are highly correlated.111 We remark that correlated hopping was used to generate effective attractions [15, 16], which may cause the pairing instability of Fermi liquid. In our work, correlated hopping directly leads to the proposed chiral superconductivity without involvement of Fermi liquid and its paring instability. Since electron hopping is complex, the two electrons exchange their place via correlated motion, and the phase factor can be arbitrary. In this case, electrons may forget their Fermi statistics. Thus, electrons may form a superconducting state via the mechanism of anyon or boson superconductivity. The above idea is very rough, but may point to a right direction [12].
@@
In this paper, we discuss a concrete realization of the above idea in 2-dimensional space. We argue that, for flat dispersion $\varepsilon\sim k^{4}$, a strong repulsive interaction may cause chiral superconductivity that spontaneously breaks time reversal and space reflection symmetry in orbital motion of electrons. Other sources of time-reversal symmetry breaking and/or spin-orbital coupling in orbital motion may further help this chiral superconductivity. Our theory is closely related to the theory of anyon superconductivity developed in Ref. 8, 12. Since a fermion is a special case of anyon, the theory of anyon superconductivity applies to fermion superconductivity with little change.
@@
Strong repulsive Coulomb interaction favor fully spin-valley polarized Fermi liquid (referred to as quarter Fermi liquid) and Wigner crystals, which have both been observed in experiments at low densities where the interaction effect is strong[17]. Furthermore, a strong coupling superconducting state that breaks time reversal and reflection symmetries in electron orbital motion was observed in Ref. 17 between quarter Fermi liquid and Wigner crystal, in tetralayer rhombohedral-stacked graphene without Moire pattern. Other superconductivities were also observed in bilayer [18, 19, 20, 21], trilayer [22, 23], and Moire-tetralayer [24] rhombohedral-stacked graphene systems.
Some BCS-pairing mechanisms for those superconducting states were explored in Ref. 25, 26, 27, 28, 29, 30, 31, 32, 33. In this paper, we take a very different approach of using only strong repulsive interaction. We find that, for flat dispersion $\varepsilon\sim k^{4}$, a repulsive Coulomb interaction may leads to chiral superconductors, which are different from BCS superconductors since they are not induced by pairing instability of Fermi liquid.
Furthermore, based on our numerical calculations, we find that, for $\varepsilon\sim k^{2}$ dispersion, the proposed chiral superconductors exhibit lower energy than the quarter Fermi liquid only at very low densities. However, at these low densities, we expect the Wigner crystal phase to have an even lower energy. In contrast, for $\varepsilon\sim k^{4}$ dispersion, the chiral superconductors demonstrate lower energy than the quarter Fermi liquid at densities near the transition to the Wigner crystal phase. This observation highlights that the $\varepsilon\sim k^{4}$ dispersion is a crucial factor in realizing chiral superconductors driven by Coulomb interactions.
@@
Some of chiral superconductors are fully spin-valley polarized or half spin-valley polarized (with two spin-valley components present and at equal density), while others are spin-valley un-polarized (with all four spin-valley components at equal density). The fully spin-valley polarized chiral superconductors are in the same phases as the spinless $p+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp$ BCS superconductor [34]. For partial spin-valley polarization, the ratio of different species of electrons are quantized as simple rational numbers. In this case, in contrast to BCS superconductors of quarter Fermi liquid, the transition from spin-valley partially-polarized chiral superconductors to quarter Fermi liquid cannot be continuous at zero temperature in the clean limit.
@@
Also, as we lower the electron density, a topological chiral superconductor is likely to change into a Wigner crystal, via a first order transition. Thus topological chiral superconductivity is more likely to appear near the transition between quarter Fermi liquid and Wigner crystal, since all those phases are driven by strong repulsive interactions. As a strong coupling superconductor, the coherence length of a chiral superconductor is about the same as the inter-electron separation.
All these chiral superconductors are topological (i.e. carry non-trivial topological order, where another example was given in Ref. 35). As a result, they carry neutral excitations with fractional self/mutual statistics and/or gapless chiral edge modes. They all break time reversal and space reflection symmetry. We expect those symmetry breaking and the associated chiral correlation (such as chiral edge state) to persist even above the superconducting transition temperature $T_{c}$. Just above $T_{c}$, the chiral superconductors should have non-zero Hall conductance of order $e^{2}/h$. Except two chiral superconductors, other chiral superconductors are not in the same phases as any BCS superconductors. For example, many of chiral superconductors have charge-4 [11] or higher condensation.
@@
Our calculation is not reliable enough to predict if a chiral superconductor can appear or not (i.e. can have energy below both quarter Fermi liquid and Wigner crystal). However, if a chiral superconductor (i.e. a superconductor that breaks time reversal and reflection symmetry) is observed in experiments near quarter Fermi liquid and Wigner crystal as in Ref. 17, our calculation suggests that it may be a topological chiral superconductor, with properties described above.
@@
The calculated ground-state energies of these topological chiral superconductors are lower, but still close to Hartree-Fock energy of a quarter Fermi liquid. To achieve an even lower energy, additional time-reversal symmetry breaking and/or spin-orbit interaction may be helpful. The strong geometric phase curvature near the bottom of the graphene band also contributes to breaking time-reversal symmetry through partial spin-valley polarization.
@@
We remark that, although most topological chiral superconductors belong to different phases than BCS superconductors, two of the topological chiral superconductors (which we denote as the $K_{2a}$-chiral superconductor and Pfaffian chiral superconductor) belong to the same phase as the “spin”-triplet or spinless $p+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp$ BCS superconductor. Here “spin” corresponds to a pair of spin-valley quantum numbers, and may not be the electron spin. Such $K_{2a}$-chiral superconductor and Pfaffian chiral superconductor can be induced by a purely repulsive interaction. They can also be induced by a pairing instability of a half Fermi liquid (for $K_{2a}$-chiral superconductor) or quarter Fermi liquid (for Pfaffian chiral superconductor) caused by an effective attractive interaction. Here, half Fermi liquid refers to a Fermi liquid formed by two species of electrons of equal density. We note that very recently, a superconducting state has been observed next to a half Fermi liquid in rhombohedral trilayer graphene in Ref. 23. The $K_{2a}$-chiral superconductor and Pfaffian chiral superconductor also have similarly low energies as other chiral superconductors at lower densities. Thus, it may also appear near the quarter Fermi liquid and Wigner crystal, as in rhombohedral tetralayer graphene in Ref. 17.
@@
II Chiral superconductivity driven by purely repulsive interactions
@@
We now examine the energetics of this state. Let $N_{\uparrow}$ ($N_{\downarrow}$) be the total number of spin-$\uparrow$ (spin-$\downarrow$) electrons. The total angular momentum of the above state is
@@
We will use (II) as a trial wave function for the associated chiral superconductor. Such a wave function is determined by $\bar{K}^{\pm}$, which must satisfy some conditions, as discussed in detail in the Appendix. In the main text, we will just summarize the results. First,
@@
$=\frac{1}{2}\sum_{I,J}N_{I}K_{IJ}N_{J}-\frac{1}{2}\sum_{I}N_{I}K_{II}$
$=\frac{N^{2}}{2}\sum_{I,J}f_{I}K_{IJ}f_{J}-\frac{N}{2}\sum_{I}f_{I}K_{II}$ (20)
where $N_{I}$ is the number of species-$I$ electrons, $N=\sum_{I}N_{I}$, and $f_{I}=\frac{N_{I}}{N}$. We remark that those $N_{I}$ electrons occupy a circular area with an area $A_{I}$. We adjust $l_{I}$ in the wave function such that $A_{I}$ for different species are all equal $A_{I}=A$.
In order to describe a superconducting or superfluid state, $K$ is required to have a single zero eigenvalue so that the total angular momentum does not contain the $N^{2}$ term:
@@
The corresponding eigenvector $f_{I}$ describes a gapless mode within the otherwise gapped chiral superconducting state. In order to be a superconducting state, we also require the electron density fluctuations giving rise to this mode to be net positive; otherwise, it would describe a charge-neutral superfluid mode instead of a superconductor. $f_{I}$ is the fraction of species-$I$ electrons. Thus, we also require
@@
There are many $K$-matrices that satisfy (21) and (22). To determine which are favorable, we first quantify the ground state energy of the chiral superconductor (II) more precisely. As described in the introduction, we consider an electron dispersion $\varepsilon=c_{2}k^{2}+c_{4}k^{4}$, where the quartic term further flattens the bottom of the band compared to a purely quadratic dispersion. From a derivation carried out in the Appendix, it can be shown that the kinetic energy per electron has the form (134)
@@
where $n_{e}$ is the total density of the electrons, and $Z_{2},\ Z_{4}$ are dimensionless parameters with the form (135)
$Z_{2}$ $=\sum_{IJ}f_{I}\bar{K}_{IJ}f_{J}+\sum_{I}f_{I}\frac{2g_{2,I}}{\pi n_{e}},$
$Z_{4}$ $=\sum_{I}2f_{I}(\sum_{J}f_{J}\bar{K}_{IJ})^{2}+\sum_{I}f_{I}\frac{16g_{4,I}}{\pi^{2}n_{e}^{2}},$
$\bar{K}$ $\equiv\bar{K}^{+}_{IJ}+\bar{K}^{-}_{IJ}.$ (26)
The dimensionless ratios, $\frac{g_{2}}{n_{e}}$ and $\frac{g_{4}}{n_{e}^{2}}$, are computed numerically via Monte Carlo method, with a current error about 10% (see Appendix).
If we assume the electrons interact via Coulomb interaction $\frac{e^{2}}{\epsilon r}$, then the interaction energy per electron is given by (142)
@@
and $g_{IJ}(z)$ is the electron pair distribution function which must be computed numerically. We find the following approximate fitting for $V_{IJ}$ (150)
@@
with error $\sim 0.03$ (see Appendix), where $\Theta(0)=0$ and $\Theta(x>0)=1$. When $\bar{K}_{IJ}=K_{\text{max}}$, the energy $E_{\text{int}}$ happen to be the same as the one component case $\bar{K}=(K_{\text{max}})$. We will use this property to compute $E_{\text{int}}$ for the case $\bar{K}_{IJ}=K_{\text{max}}$.
To determine which $K$-matrices give rise to the most stable chiral superconductors, we compute $Z_{2}$, $Z_{4}$, and $V$. The total energy per electron is given by
@@
To estimate $\bar{n}_{e}$ for the tetralayer graphene at $\gamma=4$, we use the band dispersion $10$meV $=c_{4}(0.1/a_{0})^{4}$ where $a_{0}=0.246$nm $=4.65a_{B}$, giving $\bar{n}_{e}=8.5\times 10^{12}$ cm-2. To estimate $\bar{n}_{e}$ for the tetralayer graphene at $\gamma=2$, we use band dispersion $30$meV $=c_{2}(0.1/a_{0})^{2}$, giving $\bar{n}_{e}=9.8\times 10^{12}$ cm-2.
@@
Because the densities in different spin-valley components have quantized ratios and is very different between those chiral superconducting states and quarter Fermi liquid, the zero-temperature transition between those $K$-matrix chiral superconducting states and quarter Fermi liquid is first order.